A vertical spring has a force constant of 125 N/m. By how much will the


spring be compressed if a mass of 0.525 kg is placed on top of it?

1
Expert's answer
2021-11-14T17:12:11-0500

The weight of the mass (the forces that compresses the spring) is given as follows:


"P = mg"

where "m = 0.525kg" is the mass, "g = 9.81m\/s^2" is the gravitational acceleration.

According to the Hook's law, the comression of the spring is:


"x = \dfrac{P}{k} = \dfrac{mg}{k}"

where "k = 125N\/m" is the spring constant. Thus, obtain:


"x = \dfrac{0.525kg\cdot 9.81m/s^2}{125N/m} \approx 0.0412m"

Answer. 0.0412m.